# OpenFake: An Open Dataset and Platform Toward Large-Scale Deepfake Detection

### v1

@@ 7 Conclusion
This paper introduces OpenFake, a politically grounded benchmark for evaluating deepfake detectors on high-fidelity, real-world-relevant synthetic imagery. We constructed the dataset by pairing three million politically themed real images filtered from LAION with over 963k synthetic counterparts generated using state-of-the-art open and proprietary models. To complement this static benchmark, we developed OpenFake Arena, a crowdsourced adversarial platform that incentivizes the continual creation of hard-to-detect synthetic images, enabling detectors to adapt to the evolving landscape of generative models. Through a human perception study, we showed that recent proprietary generators (e.g., Imagen 3, GPT Image 1) often fool users. We also demonstrated that detectors trained on older benchmarks fail to identify these modern fakes, while supervised models trained on OpenFake achieve strong performance, especially with compression-robust augmentations.
Our results highlight that training on the right, broad distribution is crucial. The SwinV2 model trained on OpenFake not only excels in-distribution but also transfers best to unseen generators collected from the web, outperforming baselines trained on narrower datasets. Generalization is not perfect, however; performance still varies across some proprietary or lower-quality sources, indicating that coverage and recency of training data matter.
Encouragingly, current detection methods can handle high-quality deepfakes when trained on appropriately curated and up-to-date data. OpenFake, with its real-world relevance, synthetic diversity, and extensibility through OpenFake Arena, offers a scalable path forward. By combining high-fidelity benchmarking with community-driven adversarial submissions, our framework aims to narrow the gap between generation and detection, equipping researchers and practitioners with tools to confront emerging misinformation threats in real time.
@@ Appendix A OpenFake composition and licensing
Appendix B Filtering and Captioning of LAION Images
@@
Appendix C Ethics, Privacy & Limitations
@@
Appendix D Training Details and Compute Resources
D.1 Compute Resources and Cost
@@
D.2 SwinV2 Fine-Tuning Hyperparameters
@@
D.3 Generation Parameters for Open-Source Models
@@
Appendix E OpenFake Arena
@@
Appendix F Synthetic image examples from OpenFake
[…]
@@ Appendix F Synthetic image examples from OpenFake
[…]
Figure 10: Sample images from OpenFake generated by Flux-1.1-Pro.
[…]
Figure 11: Sample images from OpenFake generated by Flux.1-Dev.
[…]

### v2

@@ 7 Conclusion
We introduced OpenFake, a politically grounded benchmark built from three million real images paired with nearly one million high-quality synthetic counterparts, and extended it with OpenFake Arena, a crowdsourced adversarial platform for continual updates. Our human perception study confirmed that recent proprietary generators often fool users, while detectors trained on older datasets fail against these models. In contrast, detectors trained on OpenFake achieved strong in-distribution performance and promising results on a curated in-the-wild set of social-media images, suggesting that reliable detection of deepfakes is attainable outside controlled benchmarks.
While performance on some proprietary or lower-quality sources remains uneven, the path forward is clear: expanding generator coverage and broadening real image diversity (e.g., camera types, capture conditions) to further improve robustness. By combining high-fidelity benchmarking with community-driven adversarial submissions, our framework aims to narrow the gap between generation and detection, equipping researchers and practitioners with tools to confront emerging misinformation threats in real time.
8 Reproducibility Statement
We have made all components necessary for reproducibility available. Section 4 describes in detail how both the real and synthetic data were collected and generated, and the complete dataset is publicly released through a permanent link (anonymous for now). The code used for training and evaluating all baseline detectors (including model weights) is provided alongside the OpenFake and in-the-wild datasets, ensuring that the reported experiments can be replicated. Appendix E gives further implementation details, including training procedures, hyperparameters, and compute resources. These resources should enable independent researchers to reproduce our results and extend the benchmarks under comparable settings.
@@ Appendix A OpenFake composition and licensing
Appendix B More results
B.1 Cross-benchmark generalization of SwinV2
Summary.
Table˜6 compares SwinV2 detectors trained on three datasets and evaluated across two external test suites. The OpenFake-trained model attains the best out-of-domain balance between TPR and TNR on both GenImage and Semi-Truths, translating into stronger Accuracy and F1. In-domain results (italicized) saturate, as expected, but are less informative about generalization.
Test set Metric Train OpenFake Train GenImage Train Semi-Truths
GenImage TPR 0.771 1.000 0.965
TNR 0.928 1.000 0.261
Accuracy 0.849 1.000 0.613
F1 Score 0.836 1.000 0.714
Semi-Truths TPR 0.909 0.830 1.000
TNR 0.962 1.000 1.000
Accuracy 0.920 0.865 1.000
F1 Score 0.947 0.907 1.000
Figure 4: t-SNE visualization of CLIP vision embeddings for 3,500 test images, including both real and synthetic images from a few generative models. Each point corresponds to an individual image, and colours indicate the generative model (or “real” for authentic images).
Why OpenFake transfers better.
We attribute the gains primarily to coverage and recency. OpenFake aggregates diverse, up-to-date generators and visual conditions (including compressions and photorealistic prompts), reducing shortcut reliance. By contrast, models trained on older or narrower distributions tend to overfit curation artifacts, which explains the high TPR but poor TNR observed when Semi-Truths-trained detectors face newer datasets: many real images are misclassified as synthetic.
Error patterns and operational trade-offs.
The cross-benchmark gaps underscore the importance of calibrating for deployment goals. A detector with inflated TPR but depressed TNR can look strong on superficially balanced metrics yet cause unacceptable false-positive rates in real pipelines. Threshold selection, confidence calibration, and cost-sensitive training are therefore critical when transferring across domains.
B.2 Robustness to compression artifacts
Real images (sourced from LAION-400M in our dataset) are typically compressed and carry authentic JPEG artifacts and blur. In contrast, synthetic images are high-resolution and minimally compressed. This mismatch creates an obvious signal: detectors may rely on compression differences instead of true semantic features, thus failing on compressed fakes. To assess this vulnerability, we implemented a data augmentation pipeline to degrade synthetic images during training, mimicking the distribution of real images, as done in previous work [16, 72, 15]. This includes random resizing, Gaussian blur, JPEG compression, and Gaussian noise. We then evaluated the SwinV2 model trained with these augmentations on a fully compressed test set and got an overall F1 score of 0.992. This demonstrates that the model remains highly accurate even when the compression signal is neutralized.
Appendix C Filtering and Captioning of LAION Images
@@
Appendix D Ethics, Privacy & Limitations
@@
Appendix E Training Details and Compute Resources
E.1 Compute Resources and Cost
@@
E.2 SwinV2 Fine-Tuning Hyperparameters
@@
E.3 Generation Parameters for Open-Source Models
@@
Appendix F OpenFake Arena
@@
Appendix G Synthetic image examples from OpenFake
@@
Figure 10: Sample images from OpenFake generated by Flux-1.1‑Pro.
Figure 11: Sample images from OpenFake generated by Flux.1‑Dev.